OpenStudy (anonymous):

(x^-2 y^3 / x^2 y)^-1/2 (x^3 y / y^1/2)?

OpenStudy (jack1):

\[(((x^{-2} y^3) / ((x^2 y))^{-1/2})) \times ((x^3 y) / (y^{1/2}))? \] \[\Huge \frac {1}{\sqrt \frac {x^{-2} y^3} {x^2 y}} \times \frac{x^3 y} { y^{1/2}} \]

OpenStudy (anonymous):

why is the first one in square root?

OpenStudy (jack1):

k this is a crazy long equation, so im gonna break it down piece at a time then build it back together so u can see where all the bitz go... \[ \large (x^-2 y^3 / x^2 y)^{-1/2} (x^3 y / y^1/2)? \] so x^-2 y^3 \[\large = \frac {y^3}{x^2} \] and x^2 y \[\large = x^2 \times y \] now to the power of half = square root ie \[\large 4^{1/2} = \sqrt 4 = 2\] and anything to a negative power = 1/ that thing to the positive power ie \[\large 4^{-1/2} = \frac {1} {\sqrt 4} = \frac 12\] following so far?

OpenStudy (anonymous):

so the 2y^2 is = 1/2 ?

OpenStudy (jack1):

\[ \large (x^{-2} y^3 / x^2 y)^{-1/2} (x^3 y / y^1/2)? \] so x^-2 y^3 \[\large = \frac {y^3}{x^2} \] and x^2 y \[\large = x^2 \times y \] so first section of equation = (x^{-2} y^3 / x^2 y)^{-1/2} \[\Huge =(\frac {\frac {y^3}{x^2}} {x^2 \times y})^{-1/2}\] \[\Huge = \frac {1}{\sqrt {(\frac {\frac {y^3}{x^2}} {x^2 \times y})}}\] with me so far?

OpenStudy (anonymous):

sorry my laptop froze

OpenStudy (anonymous):

so since its -1/2 they are all moved to the bottom?

OpenStudy (jack1):

sorry, it was getting exhausting typing in latex, gonna basic it down just to powers in plain form: first part :- x^-2 y^3 / x^2 y =x^-2 * y^3 * x^-2 * y^-1 =x^-4 * y^2 = y^2 / x^4 so (x^-2 y^3 / x^2 y)^-1/2 = ( x^-4 * y^2 )^-1/2 = ( x^2 * y^-1) = ( x^2 / y ) 2nd part: (x^3 y / y^1/2) =( x^3 * y * y^(-1/2) ) = x^3 * y^(1/2) so together 1st part times 2nd part: (x^-2 y^3 / x^2 y)^-1/2 * (x^3 y / y^1/2) = ( x^2 / y ) * x^3 * y^(1/2) = x^2 * y^-1 * x^3 * y^(1/2) = x^5 * y^-(1/2) = x^5 / y^1/2 \[ \huge \frac {x^5}{ \sqrt y } \]

OpenStudy (jack1):

so that should be the final answer, and everything above was just a manipulation of powers rules

OpenStudy (jack1):

@Ssusu

OpenStudy (anonymous):

I will look over it and make sure I got it thank you and I know this equation was long and annoying lol

OpenStudy (jack1):

sáll good ;) np